Accidents_Daily_(X) P(X=xi) 0 0.23 1 0.24 2 0.21 3 0.11 4 0.09 5 0.07 6 0.05

Compute the mean number of accidents per day.
Answer #1

We Know that the formula for mean

Mean = x*p(x)

= 0*0.23 + 1*0.24 + 2*0.21 + 3*0.11 + 4*0.09 + 5*0.07 + 6 *0.05

= 0 + 0.24 + 0.42 + 0.33 + 0.36 + 0.35 + 0.30

= 2.00

Mean = 2